# Leveraging Whisper Embeddings for Audio-based Lyrics Matching

### v1

@@ 1 Introduction
In this work, we introduce WEALY, a fully reproducible end-to-end pipeline that leverages Whisper [1] decoder embeddings, to address the abovementioned limitations and challenges, eliminating the dependency on text data or pre-existing transcriptions while providing robust and transparent baselines. This choice is further supported by recent findings which show that Whisper-based transcriptions can already provide strong results compared to annotated ground truth [2], reinforcing the value of end-to-end audio-based approaches. To evaluate our approach, we employ musical version identification (MVI) [3] as a proxy task, motivated by the challenges of data scarcity in direct lyrics matching evaluation. With it, we establish comprehensive benchmarks across three different datasets and achieve a competitive performance compared to strong baselines. We additionally conduct extensive ablation studies to analyze the impact of different losses and pooling strategies. Our ongoing work also explores multimodal extensions that integrate textual and acoustic features. We release both the code and the models’ checkpoints [^1] https://github.com/helemanc/audio-based-lyrics-matching to support transparency and reproducibility.
@@ 2 Related Work
The specific problem of lyrics matching remains largely unexplored. Prior text-based approaches [4, 5] face critical limitations, including transcription errors, spelling variations and structural inconsistencies. These approaches require extensive manual data collection and preprocessing, illustrating their limited feasibility and motivating our end-to-end audio-based method that extracts semantic lyric representations directly from audio. The lack of large-scale, high-quality lyrics datasets justifies using MVI [3] as a proxy task for large-scale evaluation, avoiding to rely on curated texts.
@@ 5 Results
Table 3: Comparison of audio content-based MVI approaches with the proposed multimodal approach the SHS dataset.

### v2

@@ 1 Introduction
In this work, we introduce WEALY (Whisper Embeddings for Audio-based LYrics matching), a fully reproducible end-to-end pipeline that leverages Whisper [1] decoder embeddings, to address the abovementioned limitations and challenges, eliminating the dependency on text data or pre-existing transcriptions while providing robust and transparent baselines. This choice is further supported by recent findings which show that Whisper-based transcriptions can already provide strong results compared to annotated ground truth [2], reinforcing the value of end-to-end audio-based approaches. To evaluate our approach, we employ musical version identification (MVI) [3] as a proxy task, motivated by the challenges of data scarcity in direct lyrics matching evaluation. With it, we establish comprehensive benchmarks across three different datasets and achieve a competitive performance compared to strong baselines. We additionally conduct extensive ablation studies to analyze the impact of different losses and pooling strategies. Our ongoing work also explores multimodal extensions that integrate textual and acoustic features. We release both the code and the models’ checkpoints [^1] https://github.com/helemanc/audio-based-lyrics-matching to support transparency and reproducibility.
@@ 2 Related Work
The specific problem of lyrics matching remains largely unexplored. Prior text-based approaches [4, 5] face critical limitations, including transcription errors, spelling variations and structural inconsistencies. These approaches require extensive manual data collection and preprocessing, illustrating their limited feasibility and motivating our end-to-end audio-based method that extracts semantic lyric representations directly from audio. The lack of large-scale, high-quality lyrics datasets justifies using MVI [3] as a proxy task for large-scale evaluation, enabling the derivation of similarity labels without relying on curated texts.
@@ 5 Results
Table 3: Comparison of audio content-based MVI approaches with the proposed multimodal approach on the SHS dataset.
@@ 6 Conclusion
7 Acknowledgments
We acknowledge ISCRA for awarding this project access to the LEONARDO supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CINECA (Italy).